OpenStudy (anonymous):

solve for y: -22=-8y+6(y-4)

OpenStudy (lgbasallote):

distribute first -22 = -8y + 6y - 24 combine like terms -22 = -2y - 24 add 24 to both sides 24 - 22 = -2y 2 = -2y divide both sides by -2 -1 = y

OpenStudy (goformit100):